OpenStudy (anonymous):

f(x)=13x^4-4x^3+4x^2+9x+14 List all zeros for function PLEASE HELP FINAL TOMORROW

OpenStudy (cwrw238):

this equation has no real roots All 4 roots are complex

OpenStudy (aum):

"List all zeros" or "List all POSSIBLE zeros" of the function? We can list all possible zeros using the rational roots theorem. But none of them will be the actual zeros.

OpenStudy (anonymous):

oh, list all possible zeros

OpenStudy (aum):

it is +/- factors of the constant term, which is 14 divided by factors of the leading coefficient, which is 13. Factors of 14: 1, 2, 7, 14 Factors of 13: 1, 13 Possible zeros:\[ \pm \frac{1, 2, 7, 14}{1,13} \]

OpenStudy (aum):

So the list of POSSIBLE factors is:\[ \pm 1, \pm\frac{1}{13}, \pm 2, \pm\frac{2}{13}, \pm 7, \pm\frac{7}{13}, \pm 14, \pm\frac{14}{13}. \]

OpenStudy (anonymous):

thank you!!

OpenStudy (aum):

you are welcome.